OpenStudy (anonymous):

What is the common difference in this sequence? 10, 17, 24, 31,...

OpenStudy (anonymous):

17-10=?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

24-17=?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

31-24=?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

so??? the answer is???

OpenStudy (anonymous):

7

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

is it hard?? NO, right? :)

OpenStudy (anonymous):

no i =m just terrible at math

OpenStudy (anonymous):

but you can do it, that 's the proof. hihihi... you CAN do everything!!